I.In each of Problems 23 through 30, use the method of reduction of order to find a second solution of the given differential equation.

1.  t^2y″ − 4ty′ + 6y = 0, y1(t) = t^2

2.  xy″ − y′ + 4x^3y = 0, x > 0; y_1(x) = sin x^2


1
Expert's answer
2022-01-05T13:41:14-0500

1.

"y_2=y_1v"

"y_2=t^2v,y_2'=2tv+t^2v',y_2''=2v+2tv'+2tv'+t^2v''=2v+4tv'+t^2v''"


"t^2(2v+4tv'+t^2v'') \u2212 4t(2tv+t^2v')+ 6t^2v = 0"

"t^2v''=0"

"v=t"

"y_2=t^3"


2.

Wronskian:

"W=y_1y_2'-y_1'y_2=ce^{-\int P(x)dx}"

where "P(x)=-1\/x"


then:

"W=cx"

"y_2'sinx^2-2y_2xcosx^2=cx"

"y_2=uv,y_2'=u'v+uv'"

"(u'v+uv')sinx^2-2uvxcosx^2=cx"


"u'v=cx\/(sinx^2)"

"v'-2vxcotx^2=0"


"dv\/v=2xcotx^2dx"

"lnv=ln(sinx^2)"

"v=sinx^2"


"du=cxdx\/(sinx^4)"

"u=\frac{4ln(sinx)+csc^2x(x(cscxcos3x-3cotx)-1)}{6}"


"y_2=\frac{4ln(sinx)+csc^2x(x(cscxcos3x-3cotx)-1)}{6}sin^2x"